If n and p are positive integers such that 8(2^p) = 4^n , what is n in terms of p?
Answer by rothauserc(4718)

8(2^p) = 4^n
:
Note rewrite above as
:
2^(3+p) = 4^n
:
now we can write
:
4^n = 2^(3+p)
:
take logarithm base 4 of both sides
:
n = ln ( 2^(3+p) ) / ln (4) for 2^p > 0 and ln is the natural logarithm
: